OpenStudy (flvs.net):

Medal And Fan The number of pages that Zak, Karen, Malia, and Ali can read in a day is shown below: Zak read 15% of his 56-page book. Karen read 12% of her 64-page book. Malia read 14% of her 68-page book. Ali read 10% of her 72-page book. Who can read the greatest number of pages in a day? Zak Karen Malia Ali

OpenStudy (flvs.net):

@jcoury

OpenStudy (flvs.net):

@Michele_Laino

OpenStudy (michele_laino):

hint: Zak reads 56*(15/100)=.. pages of his book Karen reads 64*(12/100)=...pages of her book Malia reads 68*(14/100)=...pages of her book Ali reads 72*(10/100)=...pages of his book

OpenStudy (flvs.net):

So I do 56*15/100 64*12/100 68*14/100 72*10/100?

OpenStudy (michele_laino):

yes! furthermore you have to round off each result to the nearest unit

OpenStudy (flvs.net):

So mailia is the answer right?

OpenStudy (flvs.net):

@Michele_Laino

OpenStudy (flvs.net):

@Michele_Laino

OpenStudy (michele_laino):

reassuming we have: Zak reads 56*(15/100)=8 pages Karen reads 64*(12/100)=8 pages Malia reads 68*(14/100)=10 pages Ali reads 72*(10/100)=7 pages so the answer is Malia correct!

OpenStudy (flvs.net):

YAYY!!

OpenStudy (michele_laino):

:)

OpenStudy (flvs.net):